What is the slope and y intercept of 7y + 3x +4=15 + 3x
Answer by thinkinbig2(5)

7y+3x+4=15+3x
Move all terms not containing y to the right-hand side of the equation.
7y=-3x-4+3x+15
Simplify the right-hand side of the equation.
7y=11
Divide each term in the equation by 7.
(7y)/(7)=(11)/(7)
Simplify the left-hand side of the equation by canceling the common factors.
y=(11)/(7)
To find the slope and y intercept, use the y=mx+b formula where m=slope and b is the y intercept.
y=mx+b
Using the y=mx+b formula, m=0 and b=(11)/(7).
m=0, b=(11)/(7)